OpenStudy (anonymous):

Help! will give medal. Which of the following is equivalent to |3x + 8| = 19? Select one: a. 3x + 8 = 19 or 3x + 8 = -19 b. 3x + 8 = 19 c. 3x + 8 = -19 d. none of the above

OpenStudy (anonymous):

|3x+8| = 3x+8..Absolute values are always a yield positive value. Thus |-a| = a and |a| = a

OpenStudy (anonymous):

oh duh.(: haha i keep screwing it up.

OpenStudy (anonymous):

Example: We know that 5 - 8 = -3. Well the absolute value of 5- 8 shown as |5-8| = +3.

OpenStudy (anonymous):

The answer is B